One square has a side length that is 1 footgreater than that of a smaller square. If the combined area of the two squares is 113 feet. find the length of the side of the smaller square.
Answer by fractalier(6550)

Call the side of the smaller square, x.
Thus the larger square has a side of x+1.
Since area is side squared, we have
x^2 + (x+1)^2 = 113
x^2 + x^2 + 2x + 1 = 113
2x^2 + 2x - 112 = 0
x^2 + x - 56 = 0
(x + 8)(x - 7) = 0
X = -8 or x = 7
Can't be negative so you have a side of
7 feet